If cosx=tany, cosy=Tanz and cosz= tanx, prove that |sinx|=|siny|=|sinz|

Grade:11

1 Answers

Vikas TU
7 years ago
convert evrythng in cos form.
cosx  = cos(1/root(1+y^2))
x = 1/root(1+y^2) …......................................(1)
 
cosy =  cos(1/root(1+z^2))
y =  1/root(1+z^2) …......................................(2)
 
similarly,
 
cosz =  cos(1/root(1+x^2))
z =  1/root(1+x^2)................................(3)
 
Take sinboth sides in all 3 eqns. and compare,
and take mod, thus the magnitude are all in the form of square form which are all same